(#9) use the laplace transform to solve to given differential equation to the indicated initial conditions. where appropriate, write 'f' in terms of unit step functions.
8. y-4y 0, y(0) = 0 = 0. v(0) = 4 9. y-4y+4y te2, y(0) 1
Answer #1

9 Y - 4y + 4y = te2+ 410) - 0,460) = 0 Taking laplace both side te {41-441 +44 3 = 2m (+3 6 243 en 44} - 4 h {y) +4 2942 -